I have two cars in my drive way. one is red and the other is green.


a) the probability that the red one is out of gas is .7. given the red one is out of gas, the probability the green one is out of gas is .6. what is the probability that both are out of gas?

b) the probablity that both cars need to be washed is .5. the probability that the red on needs washing is .7. what is the probability that the green one needs washed given the red car needs washing?
1
Expert's answer
2011-10-11T08:41:40-0400
a) a. P=0.7*0.6=0.42
b) b. P=0.5+0.7-0.5*0.7=0.85